# MAPLE: Enhancing Review Generation with Multi-Aspect Prompt LEarning in Explainable Recommendation

### v1

@@ 4.1 Dataset
Unlike hotels with limited differentiating factors (e.g., bed comfort, room size), different restaurants offer a wide variety of unique items (e.g., specific dishes like tonkatsu ramen for a Japanese ramen store; risotto for an Italian restaurant). This distinctiveness in aspect terms allows for more precise and compelling explanations for a specific restaurant. Therefore, we focus on the restaurant domain by benchmarking against two datasets from Yelp, which we refer to as Yelp19 Li et al. (2020) and Yelp23 (Inc., 2023), distinguishing them based on the year the data was provided. We provide the dataset sources, statistics and the details of how we renew the datasets using the automated sentiment-pipeline in Appendix A.
@@
Factuality Aspect-wise Exp. Text Diversity
Method iFMR FCR iFCR GT-FMR USR uUSR iUSR D-2 D-3 ENTR
Att2seq 0.693 0.055 0.063 0.082 0.348 0.993 0.864 0.802 0.858 8.389
NRT 0.672 0.051 0.063 0.075 0.384 0.993 0.894 0.804 0.867 8.331
PETER 0.704 0.062 0.057 0.096 0.263 0.991 0.733 0.819 0.858 8.173
PEPLER 0.661 0.047 0.047 0.087 0.301 0.991 0.797 0.365 0.412 8.241
MAPLE 0.807 0.185 0.108 0.087 0.951 0.999 0.997 0.684 0.808 11.015
MAPLE-GT 0.684 0.148 0.086 0.167 0.475 0.994 0.890 0.801 0.867 9.362
@@
Factuality Aspect-wise Exp. Text Diversity
Method iFMR FCR iFCR GT-FMR USR uUSR iUSR D-2 D-3 ENTR
Att2seq 0.618 0.012 0.022 0.230 0.253 0.987 0.977 0.878 0.851 8.024
NRT 0.615 0.009 0.021 0.224 0.216 0.975 0.956 0.876 0.850 7.657
PETER 0.678 0.015 0.024 0.251 0.272 0.962 0.934 0.867 0.847 8.022
PEPLER 0.690 0.013 0.025 0.261 0.303 0.950 0.911 0.809 0.799 8.311
MAPLE 0.794 0.063 0.041 0.206 0.855 0.999 0.997 0.816 0.866 11.485
MAPLE-GT 0.691 0.039 0.029 0.333 0.620 0.998 0.994 0.892 0.886 10.316
Table 1: Automatic evaluation results on 10,000 user-item pairs sampled from the test sets. The "i"-prefix is "item-wise" for short and "u"-prefix is "user-wise". Best and second-best performances are bold-faced and underlined, respectively. Since MAPLE-GT employs ground-truth aspects, we do not put this variant into comparisons. Unless otherwise stated, MAPLE in the paper defaults to the Supervised@3 aspect-recommendation strategy.
@@ 4.2.1 Factuality
Factuality metrics assess how factual the model’s generated content is. Xie et al. (2023) introduces the Entailment Ratio, which measures how many statements are supported by existing reviews of the same item. However, the generated content can synthesize aspects and opinions from multiple reviews, which the Entailment Ratio metric may incorrectly classify as "unfactual." To address the above issue, We propose Item-wise Feature Matching Ratio (iFMR). We treat the collection of extracted features [^2] We use ”features” interchangeably with ”aspect terms”, because Li et al. (2020) refer to aspect terms as features and use it in naming the proposed automatic metrics. in one restaurant’s reviews as its "menu"; our rationale is that if the generated content contains any feature from the menu, it is "factual." Formally,
@@
4.2.2 Aspect-wise Explainability
To assess the aspect-wise diversity and coverage, we adopt two metrics from Li et al. (2020): Feature Coverage Ratio (FCR) and Ground-Truth Feature Matching Ratio (GT-FMR). FCR is computed as the number of distinct features contained in all the generated explanations, divided by the total number of the features in the training set  [^3] Because we only evaluate a subset of testset, we collect the features associated with the items in the subset of testset, and use its total count as denominator.. Ground-Truth Feature Matching Ratio (GT-FMR) measures whether a generated explanation contains the feature in the ground-truth text.  [^4] It was originally proposed as Feature-Matching Ratio (FMR). To distinguish it from the factuality metric item-wise FMR, we prefix it with ”GT”. Additionally, it’s important to consider the perspective of restaurant owners, who invariably prefer that a greater number of their restaurant’s advantages are highlighted and recommended to users in the explanation texts. Therefore, we craft item-wise Feature Coverage Ratio (iFCR) to provide insights into how well a model covers the relevant features for each item based on the training data. For more implementation details and concrete formulas, refer to Appendix B.1.
4.2.3 Informativeness
We measure the informativeness of generated sentences using token, sentence, and corpus level evaluations. For token-based merics, we use the Distinct-N metrics by Li et al. (2016). To assess the severity of repetitive sentences generated by the model, we utilize the Unique-Sentence Ratio (USR) Li et al. (2020) and the derived metrics, user-wise USR and item-wise USR [^5] User-wise USR is evaluated from a user’s perspective by calculating the USR ratio within their respective collections of generated explanations, and then averaging the results. Same for item-wise USR.. For corpus-level, we employ entropy-based metric ENTR by Jhamtani et al. (2018).
4.3 Comparison Methods
We compare our model against two groups of baseline methods. The first group consists of classic end-to-end review generation models focused on NLG for explanations. We cover Att2seq (Dong et al., 2017), NRT (Li et al., 2017), PETER (Li et al., 2021), PEPLER (Li et al., 2023b). The second group is the retriever architecture in the retriever-reader framework, in which to the best of our knowledge only the personalized retriever in PRAG (Xie et al., 2023) is available. For more details, see Appendix B.2.
@@ 5.1 Quantitative Analysis on Explanations
In our quantitative analysis, we compare MAPLE with baseline review-generation models that use item and user IDs but lack aspect input signals. Despite not multitasking on overall rating prediction, MAPLE significantly outperforms in factuality, feature coverage, and both sentence-level and corpus-level diversity. It shows more than a 10% improvement in item-wise FMR, recommending accurate features in about 80% of explanations. While striving for precision, such as choosing specific features like "salmon sushi" over general "food," MAPLE experiences a trade-off between FCR and FMR, especially notable in its lower GT-FMR score on Yelp23. Nonetheless, it excels in textual diversity on Yelp19, evidenced by high USR and ENTR scores, indicating minimal sentence repetition and high creativity. For qualitative analysis and case studies, refer to Appendix D.
@@ 5.2 Ablation Study on Aspect Recommendation Strategies
Factuality Aspect-wise Exp. Text Diversity Text Sim. Ranking
Method iFMR iFCR GT-FMR USR D-2 ENTR BLEU-4 HR
S@1 0.728 0.077 0.061 0.654 0.830 10.580 0.297 0.350
S@2 0.789 0.099 0.081 0.840 0.768 11.094 0.245 0.568
S@3 0.805 0.108 0.087 0.944 0.721 10.958 0.238 0.716
S@4 0.823 0.125 0.096 0.971 0.666 11.497 0.182 0.806
Heuristic@3 0.731 0.094 0.076 0.942 0.733 10.998 0.230 0.419
GT@1 0.684 0.086 0.167 0.728 0.833 8.041 0.585 -
@@
S@1 0.548 0.023 0.232 0.336 0.877 8.587 1.528 0.710
S@2 0.725 0.035 0.200 0.661 0.856 11.487 0.952 0.863
S@3 0.794 0.041 0.207 0.855 0.816 11.485 0.947 0.916
S@4 0.834 0.054 0.227 0.921 0.734 12.182 0.596 0.948
Heuristic@3 0.785 0.042 0.178 0.855 0.816 11.441 0.861 0.811
GT@1 0.691 0.029 0.333 0.620 0.892 10.316 2.616 -
Table 2: Comparison of MAPLE with different aspect selection strategies. S@k denotes Supervised strategies with different number of $K$. HR is short for Hit Ratio. For Yelp23, there might be multiple ground-truth review segments, in this case we report its BLEU-4 with multi-reference mechanism.
@@
    Supervised@K (K=1,2,3): Inferencing the trained aspect-recommendation component to select top aspects from a trimmed predicted aspect distribution (Appendix C.1).
@@
    GT@1: Simulate the scenario where the user has either specified their interested aspect or the aspect-recommendation model is 100% accurate by directly using the ground-truth aspect.  [^7] We do not inference the trained aspect-recommendation component for Heuristic and GT.
A proper mix of aspects enriches explanations. Increasing the $K$ value in the model enhances the hit ratio, which boosts corpus-level diversity and feature coverage. Conversely, it slightly reduces token-level diversity, as evidenced by lower Distinct-2 scores, and significantly diminishes text-similarity scores. This reduction in Distinct-2 scores is attributed to the lengthening of sentences as $K$ increases. Given that the BLEU (Papineni et al., 2002) is considered an indicator to textual relevance to the ground-truth reviews, we attempt to strike a balance in between the diversity and textual relevance by monitoring BLEU-4 score. We employ the elbow method on BLEU-4 to identify the point where the slope sharply changes (when the component starts to produce overly generalized selection) and choose the point as our optimal $K$. Additionally, while heuristic selection strategies yield suboptimal results in aspect-wise explainability, the supervised approach is preferred due to its higher aspect ranking scores.
@@
Ground-truth aspects lead to precise feature prediction and higher textual relevance. When provided with the ground-truth aspect category, the model’s GT-FMR score significantly improves, demonstrating effective training of aspect prompts. This improvement is attributed to the accurate aspect signal, which narrows the feature pool and thereby enhances feature prediction accuracy. Furthermore, when conditioning on the ground-truth aspect, the BLEU-4 score tops all other strategies, showing that under the scenario where the user is able to provide the aspect category he/she is interested in, the explanation precision significantly improves. This underscores the value of precise user input in enhancing the relevance and accuracy of generated explanations.
@@ 5.3 Explainability of Aspect Prompts
Our hypothesis posits that MAPLE’s aspect prompt tokens better retain and utilize less frequent features, which we explored using t-Distributed Stochastic Neighbor Embedding (t-SNE) (van der Maaten and Hinton, 2008) to visualize the semantic clustering of feature words and aspect prompts. Figure 3 shows linear clusters for the aspects and their feature words, confirming MAPLE’s efficient learning of semantic associations. We further calculate the FCR scores by only considering the head 5 and the tail 5 aspect categories, denoted head FCR and tail FCR, and then present the statistics in Table 3. These results suggest that MAPLE’s aspect prompt tokens significantly improve its ability to remember and utilize even the less frequent features, maintaining competitive performance even for less common aspect categories.
@@ 5.4 MAPLE as a Discrete Retriever
Latent Perspective. We project MAPLE-generated explanations into the same latent space as PRAG’s queries by encoding it with all-mpnet-base-v2 to assess similarity against the ground-truth review, using Cosine Similiarity and Mean Squared Error (MSE). Since PRAG’s objective during training is to minimize the MSE, it performs very well in this context.
Aspect Perspective. We focus on the models’ ability to identify relevant aspects. MAPLE employs an aspect recommendation module that targets relevant losses, whereas for PRAG, we utilize its retrieved reviews from the user and item history, and collect the top-3 reviews’ corresponding aspects as its prediction. Although MAPLE’s queries show only a 30% similarity to the ground-truth reviews based on latent metrics, it more effectively identifies the correct aspects, reflecting a stronger alignment with the target aspect goals. Overall, PRAG excels in generating queries resembling the ground-truth reviews and MAPLE shows superior performance in accurately targeting the relevant aspects.
@@ 7 Limitations
Despite its high feature and textual diversity, the MAPLE model presents several limitations. A notable challenge is the labeling of aspect categories. Although automated, this process still necessitates manual effort to define the aspect category inventory. The quality of these labels and their distribution across the dataset impact the training difficulty of the aspect recommendation component and subsequently the inferenced text quality and style. In cases where the label distribution is highly skewed, it might be necessary to optimize the aspect recommendation component separately. Additionally, our introduction of the item-wise Feature Matching Ratio marks a pioneering step towards enriching the aspect-wise factuality perspective of explainable recommendation model evaluations. While this metric adeptly identifies factual features in sentences, its capability to detect non-factual elements still needs future research and methodological advancement.

### v2

@@ 4.1 Dataset
In our experiments, we focus on the restaurant domain by benchmarking against two datasets from Yelp, which we refer to as Yelp19 Li et al. (2020) and Yelp23 (Inc., 2023), distinguishing them based on the year the data was provided. Different restaurants offer a wide variety of unique items (e.g., specific dishes like tonkatsu ramen for a Japanese ramen store; risotto for an Italian restaurant). This distinctiveness in aspect terms allows for more precise and compelling explanations for a specific restaurant. We provide the dataset sources, statistics, and details of how we renew the datasets using the automated sentiment pipeline in Appendix A.
@@
Factuality Aspect-wise Exp. Text Diversity Gen. Quality
Method iFMR FCR iFCR GT-FMR USR uUSR iUSR D-2 D-3 ENTR MAUVE
Att2seq 0.693 0.055 0.063 0.082 0.348 0.993 0.864 0.802 0.858 8.389 0.0404
NRT 0.672 0.051 0.063 0.075 0.384 0.993 0.894 0.804 0.867 8.331 0.0402
PETER 0.704 0.062 0.057 0.096 0.263 0.991 0.733 0.819 0.858 8.173 0.0239
PEPLER 0.661 0.047 0.047 0.087 0.301 0.991 0.797 0.365 0.412 8.241 0.0061
ERRA 0.775 0.060 0.064 0.097 0.277 0.917 0.859 0.825 0.868 8.316 0.0267
MAPLE 0.807 0.185 0.108 0.087 0.951 0.999 0.997 0.684 0.808 11.015 0.0699
[…]
@@
Factuality Aspect-wise Exp. Text Diversity Gen. Quality
Method iFMR FCR iFCR GT-FMR USR uUSR iUSR D-2 D-3 ENTR MAUVE
Att2seq 0.618 0.012 0.022 0.230 0.253 0.987 0.977 0.878 0.851 8.024 0.0634
NRT 0.615 0.009 0.021 0.224 0.216 0.975 0.956 0.876 0.850 7.657 0.0486
PETER 0.678 0.015 0.024 0.251 0.272 0.962 0.934 0.867 0.847 8.022 0.0604
PEPLER 0.690 0.013 0.025 0.261 0.303 0.950 0.911 0.809 0.799 8.311 0.0426
ERRA 0.775 0.033 0.03 0.181 0.388 0.963 0.858 0.884 0.878 8.992 0.0267
MAPLE 0.794 0.063 0.041 0.206 0.855 0.999 0.997 0.816 0.866 11.485 0.2183
[…]
Table 1: Automatic evaluation results on 10,000 user-item pairs sampled from the test sets. The "i"-prefix is "item-wise" for short, and the "u"-prefix is "user-wise". The best and second-best performances are bold-faced and underlined, respectively. Since MAPLE-GT employs ground-truth aspects, we do not put this variant into comparisons. Unless otherwise stated, MAPLE in the paper defaults to the Supervised@3 aspect-recommendation strategy.
@@ 4.2 Automatic Evaluation Metrics
4.2.1 Diversity
We measure the diversity of generated sentences using token, sentence, and corpus-level evaluations. For token-based metrics, we use the Distinct-N metrics by Li et al. (2016). To assess the severity of repetitive sentences generated by the model, we utilize the Unique-Sentence Ratio (USR) Li et al. (2020) and the derived metrics, user-wise USR and item-wise USR [^3] User-wise USR is calculated as the average USR ratio across a user’s generated explanations, with the same approach applied for item-wise USR.. For the corpus-level, we employ the entropy-based metric ENTR by Jhamtani et al. (2018).
@@ 4.2.2 Factuality
For factuality evaluation, Xie et al. (2023) introduces the Entailment Ratio, which measures how many statements are supported by existing reviews of the same item. However, the generated content can synthesize aspects and opinions from multiple reviews, which the Entailment Ratio metric may incorrectly classify as "unfactual." To address the above issue, we propose Item-wise Feature Matching Ratio (iFMR). We treat the collection of extracted features [^4] We use ”features” interchangeably with ”aspect terms” as Li et al. (2020) refer to aspect terms as features in their proposed metrics. in one restaurant’s reviews as its "menu"; our rationale is that if the generated content contains any feature from the menu, it is "factual." Formally,
@@
4.2.3 Precision
To evaluate precision, which can also be interpreted as aspect-wise explainability, we adopt two metrics from Li et al. (2020): Feature Coverage Ratio (FCR) and Ground-Truth Feature Matching Ratio (GT-FMR). FCR is computed as the number of distinct features contained in all the generated explanations, divided by the total number of the features in the training set [^5] Since we evaluate only a test subset, we count the features associated with its items and use the total as the denominator.. Ground-Truth Feature Matching Ratio (GT-FMR) measures whether a generated explanation contains the feature in the ground-truth text. [^6] It was originally proposed as a Feature-Matching Ratio (FMR). To distinguish it from the factuality metric item-wise FMR, we prefix it with ”GT”. Additionally, it’s important to consider the perspective of restaurant owners, who invariably prefer that a greater number of their restaurant’s advantages be highlighted and recommended to users in the explanation texts. Therefore, we craft item-wise Feature Coverage Ratio (iFCR) to provide insights into how well a model covers the relevant features for each item based on the training data. For more implementation details and concrete formulas, refer to Appendix B.1.
4.2.4 Generation Quality
Unlike targeted generation tasks such as text summarization, explanation generation is a form of open-ended text generation. This nature makes the use of metrics like BLEU or ROUGE, which rely on exact token-by-token matches, less suitable. To address this, we employ MAUVE Althubyani et al. (2024) to quantify the gap between generated explanations and human texts, providing a statistical measure of generation quality.
4.3 Baselines
We compare our model against two groups of baseline methods. The first group consists of classic end-to-end review generation models focused on NLG for explanations. We cover Att2seq (Dong et al., 2017), NRT (Li et al., 2017), PETER (Li et al., 2021), PEPLER (Li et al., 2023), and ERRA (Cheng et al., 2023). The second group is the retriever architecture in the retriever-reader framework, in which, to the best of our knowledge, only the personalized retriever in PRAG (Xie et al., 2023) is available. For more details, see Appendix B.2.
@@ 5.1 Quantitative Analysis on Explanations
In our quantitative analysis, we compare MAPLE with baseline review-generation models that use item and user IDs but lack aspect input signals. Despite not multitasking on overall rating prediction, MAPLE significantly outperforms in factuality, feature coverage, and both sentence-level and corpus-level diversity. It shows more than a 10% improvement in item-wise FMR, recommending accurate features in about 80% of explanations. While striving for precision, such as choosing specific features like "salmon sushi" over general "food," MAPLE experiences a trade-off between FCR and FMR, especially notable in its lower GT-FMR score on Yelp23. Nonetheless, it excels in textual diversity on Yelp19, evidenced by high USR and ENTR scores, indicating minimal sentence repetition and high creativity. Besides, MAPLE also achieved high MAUVE scores, demonstrating a generation distribution closely aligned with human texts. For qualitative analysis and case studies, refer to Appendix D.
@@ 5.2 Ablation Study on Aspect Recommendation Strategies
Factuality Aspect-wise Exp. Text Diversity Gen. Qty. Text Sim. Ranking
Method iFMR iFCR GT-FMR USR D-2 ENTR MAUVE BLEU-4 HR
S@1 0.728 0.077 0.061 0.654 0.830 10.580 0.0560 0.297 0.350
S@2 0.789 0.099 0.081 0.840 0.768 11.094 0.0451 0.245 0.568
S@3 0.805 0.108 0.087 0.944 0.721 10.958 0.0699 0.238 0.716
S@4 0.823 0.125 0.096 0.971 0.666 11.497 0.0813 0.182 0.806
Heuristic@3 0.731 0.094 0.076 0.942 0.733 10.998 0.0504 0.230 0.419
[…]
@@
S@1 0.548 0.023 0.232 0.336 0.877 8.587 0.0548 1.528 0.710
S@2 0.725 0.035 0.200 0.661 0.856 11.487 0.2018 0.952 0.863
S@3 0.794 0.041 0.207 0.855 0.816 11.485 0.2183 0.947 0.916
S@4 0.834 0.054 0.227 0.921 0.734 12.182 0.1594 0.596 0.948
Heuristic@3 0.785 0.042 0.178 0.855 0.816 11.441 0.1796 0.861 0.811
[…]
Table 2: Comparison of MAPLE with different aspect selection strategies. S@k denotes Supervised strategies with different values of $K$. HR is short for Hit Ratio. For Yelp23, there might be multiple ground-truth review segments; in this case, we report its BLEU-4 with the multi-reference mechanism.
@@
    Supervised@K (K=1,2,3): Inferencing the trained aspect-recommendation component to select $K$ aspects from a trimmed predicted aspect distribution (Appendix C.1).
@@
    GT@1: Simulating the scenario where the user has either specified their interested aspect or the aspect-recommendation model is 100% accurate by directly using the ground-truth aspect. [^8] We do not inference the trained aspect-recommendation component for Heuristic and GT.
A proper mix of aspects enriches explanations. Increasing the $K$ value in the model enhances the hit ratio, which boosts corpus-level diversity and feature coverage. Conversely, it slightly reduces token-level diversity, as evidenced by lower Distinct-2 scores, and significantly diminishes text-similarity scores. This reduction in Distinct-2 scores is attributed to the lengthening of sentences as $K$ increases. Given that the BLEU (Papineni et al., 2002) is considered an indicator of textual relevance to the ground-truth reviews, we attempt to strike a balance between the diversity and textual relevance by monitoring BLEU-4 score. We employ the elbow method on BLEU-4 to identify the point where the slope sharply changes (when the component starts to produce overly generalized selection) and choose the point as our optimal $K$. We choose $K=3$ as the default value. Additionally, while heuristic selection strategies yield suboptimal results in aspect-wise explainability, the supervised approach is preferred due to its higher aspect ranking scores.
@@
Ground-truth aspects lead to precise feature prediction and higher textual relevance. When provided with the ground-truth aspect category, the model’s GT-FMR score significantly improves, demonstrating effective training of aspect prompts. This improvement is attributed to the accurate aspect signal, which narrows the feature pool and thereby enhances feature prediction accuracy. Furthermore, when conditioning on the ground-truth aspect, the BLEU-4 score tops all other strategies, showing that under the scenario where the user can provide the aspect category he/she is interested in, the explanation precision significantly improves. This underscores the value of precise user input in enhancing the relevance and accuracy of generated explanations.
@@ 5.3 Explainability of Aspect Prompts
Our hypothesis posits that MAPLE’s aspect prompts tokens better retain and utilize less frequent features, which we explored using t-Distributed Stochastic Neighbor Embedding (t-SNE) (van der Maaten and Hinton, 2008) to visualize the semantic clustering of feature words and aspect prompts. Figure 4 shows linear clusters for the aspects and their feature words, confirming MAPLE’s efficient learning of semantic associations. We further calculate the FCR scores by only considering the head 5 and the tail 5 aspect categories, denoted head FCR and tail FCR, and then present the statistics in Table 3. These results suggest that MAPLE’s aspect prompt tokens significantly improve its ability to remember and utilize even the less frequent features, maintaining competitive performance even for less common aspect categories.
@@ 5.4 MAPLE as a Discrete Retriever
Latent Perspective. We project MAPLE-generated explanations into the same latent space as PRAG’s queries by encoding it with all-mpnet-base-v2 to assess similarity against the ground-truth review, using Cosine Similarity and Mean Squared Error (MSE). Since PRAG’s objective during training is to minimize the MSE, it performs very well in this context.
Aspect Perspective. We evaluate the model’s ability to identify relevant aspects. MAPLE uses an aspect recommendation module targeting relevant losses, whereas for PRAG, we utilize its retrieved reviews from the user and item history, and collect the top-3 reviews’ corresponding aspects as its prediction. Despite MAPLE’s queries showing only 30% similarity to ground-truth reviews, it more effectively identifies correct aspects, demonstrating stronger alignment with target goals. Overall, PRAG excels in generating queries similar to ground-truth reviews, while MAPLE outperforms in accurately targeting relevant aspects.
@@ 7 Limitations
Despite its high features and textual diversity, the MAPLE model presents several limitations. A notable challenge is the labeling of aspect categories. Although automated, this process still necessitates manual effort to define the aspect category inventory. The quality of these labels and their distribution across the dataset impact the training difficulty of the aspect recommendation component and subsequently the inference text quality and style. In cases where the label distribution is highly skewed, it might be necessary to optimize the aspect recommendation component separately. Additionally, our introduction of the item-wise Feature Matching Ratio marks a pioneering step towards enriching the aspect-wise factuality perspective of explainable recommendation model evaluations. While this metric adeptly identifies factual features in sentences, its capability to detect non-factual elements still needs future research and methodological advancement.
Acknowledgments
We thank the Google Pixel Software Team for providing strong research support to this project.